OpenStudy (anonymous):

Determine whether the problem should be solved using the formula for an arithmetic sequence, arithmetic series, geometric sequence, or geometric series. Explain your answer in complete sentences. You do not need to solve. Jackie deposited $5 into a checking account in February. For each month following, the deposit amount was doubled. How much money was deposited in the checking account in the month of August?

OpenStudy (anonymous):

@snowman97 Can you help me out ?:)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Do you know the answer to my problem?

OpenStudy (anonymous):

have answer choices

OpenStudy (anonymous):

Determine whether the problem should be solved using the formula for an arithmetic sequence, arithmetic series, geometric sequence, or geometric series. Explain your answer in complete sentences. You do not need to solve.

OpenStudy (anonymous):

you have to use gp with r=5, and then sum i hope this helps

OpenStudy (anonymous):

@yashdhanuka Thank you so much, can you help me with a few more of these exact problems.. they confuse me

OpenStudy (anonymous):

A local grocery store stacks the soup cans in such a way that each row has 2 fewer cans than the row below it. If there are 32 cans on the bottom row, how many total cans are on the bottom 14 rows?

OpenStudy (anonymous):

yes sure:D will be glad to help you.

OpenStudy (anonymous):

shouldn't the geometric series have common ratio as 2???

OpenStudy (anonymous):

Wait what does gp stand for haha

OpenStudy (anonymous):

omg! am sorry yes i typed 5 by mistaske xavier is absolutely right.

OpenStudy (anonymous):

haha but what does gp stand for?

OpenStudy (anonymous):

geometric progression :D

OpenStudy (anonymous):

so it is geometric series?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so A local grocery store stacks the soup cans in such a way that each row has 2 fewer cans than the row below it. If there are 32 cans on the bottom row, how many total cans are on the bottom 14 rows? this is geometric series?

OpenStudy (anonymous):

This is a question of Arithmetic Progression and not that of a GP. Let number of cans in the top most row be = a In this case the common difference (d) = 2 According to the question there are 14 rows in all. And the Bottom Row ( 14th ) has 32 cans. Hence ----- => 32 = [ a + ( 14 - 1 ) 2 ] => a = 6 We have now to calculate total number (N) of cans in the whole lot. => N = ( 14/2 ) [ 6 + 32 ] = 266 .

OpenStudy (anonymous):

nope... here your number of cans in each row always differ by 2... i.e. d,d-2,n-4,n-6 and so on... this is an arithmetic series

OpenStudy (anonymous):

@nomad96 did you understand?

OpenStudy (anonymous):

Alright now what would Jackie deposited $5 into a checking account in February. For each month following, the deposit amount was doubled. How much money was deposited in the checking account in the month of August? be

OpenStudy (anonymous):

nomad until and unless you do the sums on your own you will never understand :D

OpenStudy (anonymous):

but it says you dont have to solve them

OpenStudy (anonymous):

I already know how to solve sequences and series, I did 4 lessons on it

OpenStudy (anonymous):

:) thats really nice :D

OpenStudy (anonymous):

So can you just help me? I really am stuck on this

OpenStudy (anonymous):

sure a[r^n-1]/[r-1] so by substituion as there are 6 months n=6 so we get 5 * 63 :D

OpenStudy (anonymous):

and what sequence or series would that be?

OpenStudy (anonymous):

its an gp and we use the formula for sum of gp :D

OpenStudy (anonymous):

geometric sequence?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

yay thanks :) what about A major US city reports a 12% increase in decoration sales during the yearly holiday season. If decoration sales were 8 million in 1998, how much did the city report in total decoration sales by the end of 2004?

OpenStudy (anonymous):

Total no of years to be considered = ( 2004 - 1998 ) = 6 years. According to the problem, Total sale in the year 1998 was = 8 M , Hence -- Total sale in 2004 => 8 [ 1 + (12/100) ]⁶ = 15.79058148 M

OpenStudy (anonymous):

So it would be geometric series

OpenStudy (anonymous):

yes :D

OpenStudy (anonymous):

A fireplace contains 46 bricks along its bottom row. If each row above decreases by 4 bricks, how many bricks are on the 12th row? This would be arithmetic sequence?

OpenStudy (anonymous):

yes :) whats your name?

OpenStudy (anonymous):

nomad

OpenStudy (anonymous):

sweet name <3

OpenStudy (anonymous):

I was just kidding but thanks

OpenStudy (anonymous):

Using complete sentences, explain the difference between an exponential function and a geometric series.

OpenStudy (anonymous):

last one !!:)

OpenStudy (anonymous):

Let r = e^(1/n) Integral_[0,1] e*x dx = lim_(n --> oo) [1 + r + r^2 + ... + r^(n-1)]/n = ( r^n - 1)/ n ( r - 1) = (e - 1)/ denominator denominator D(n) = (e^(1/n) - 1)/(1/n) --> d/dx (e^(x)) at x = 0 = 1 Therefore Integral _[0.1] e^x dx = e - 1 ---- Interest : rate r %compounded every 3 months. Money in account after 1 year: M)1) = (1 + r/400)^4 M(0) Compound n times per year :M(1) = ( 1 + (1/n)(r/100))^n M(0)) Continuous compounding: n --> oo M(1) = e^(r/00) M(0) Best I can do.

OpenStudy (anonymous):

why dont you reply? :(

OpenStudy (anonymous):

Sorry I stepped outside for a min @yashdhanuka